OpenStudy (anonymous):

write the quation of a hyperbola with the center at the orgin having verticies (0,-7) and (0,7) and the foci at (0,-9) and (0,-9)

OpenStudy (anonymous):

thats not an equation of a hyperbola...

OpenStudy (anonymous):

hold on... gimme a minute

OpenStudy (anonymous):

it is y^2/49 -x^2/32 =1

OpenStudy (anonymous):

how do you figure that out?

OpenStudy (anonymous):

a = distance of vertex from center. so a=7 c = distance of focus from center =9 b^2 = c^2-a^2 = 81-49=32 since the vertices are (0,-7) and (0,7), it is symmetrical across the x axis so it has the equation y^2/a^2 - x^2/b^2 =1